OpenStudy (anonymous):

Find the magnitude and direction angle θ of the following vectors. Round the magnitude to the nearest tenth, and round the direction angle to the nearest degree, if rounding is necessary. v = (-5, -2) Choice 'A' |v| ≈ 5.4; θ ≈ 22° Choice 'B' |v| ≈ 5.4; θ ≈ -22° Choice 'C' |v| ≈ 26.4; θ ≈ 7° Choice 'D' |v| ≈ 5.4; θ ≈ 12°

ganeshie8 (ganeshie8):

\(\large \vec{v}= (x, y)\) \(\large |\vec{v}| = \sqrt{x^2+y^2}\) \(\large \theta = \tan^{-1}(\frac{y}{x})\)